Instructions for Changing PoPs During the Night
It is proper protocol for the PI's, remote or visiting scientists to announce any PoP changes 10 minutes before needed so the operator can bring up the required windows and look up the values for the change. Please give the PoP's for only the telescopes that need changing. Use this format: E1 3-> 1, E2 3-> 5, W1 5-> 3, etc. Give the current PoP and the desired PoP only.
  • Unlock Tiptilt on all beams using the Tiptilt Servo GUI.
  • If the telescopes are still pointing at a star, make sure to periodically check the AZ/EL status on the telescope servers during the POP change procedure. If the telescopes ever get too low or start drifting, click [STOW] to safely stow the telescopes until after the POP change finishes.
  • Make sure that the cart involved is further back than the home sensor.
  • Close the M5 cover on each telescope that needs a POP change using the telescope gui Control Tab.
  • Make sure the telescope TVs are on and click [M10AL] on the Main Tab on each telescope GUI.
  • Under ESP Controllers menu, open "Periscopes S1/S2/W1" and "Periscopes W2/E2/E1"
  • Under Pico Controllers menu, open "Pico #2"
  • Open the POPs GUI. Click [Overview] to bring up the overview to see the status of which POPs are in.
    • When reading the overview, the light starts from the top and goes down. The light will reflect off the first POP that is in.
    • "--" means the POP mirror is in place, "|" means the POP mirror is out, "N" means the POP mirror is moving or has not hit the limit switch.
    • If the POP overview reports nonsense, then click "FLUSH" on the PoPS GUI, followed by "REOPEN" on the PoPS GUI, and finally, "UPDATE" on the POP Overview. The overview should now report the correct PoPs.
  • Send VisBeams to the correct pair of telescopes that you need to adjust.
  • Open the "B1" through "B6" shutters on the SHUTTERS GUI for the telescopes that you need to adjust. Also open the [LASER] shutter on the SHUTTERS GUI.
  • Turn the Laser Filter to "ND 0.0".
  • Check to make sure the counts are not saturating on the TT server (<200,000). If the counts are above 200,000 it might be a good idea to go into the lab and put the TT cover on until after the POP changes are finished. Remember to wear goggles if the metrology laser is on.
  • You should now see the laser appear on the "M10AL" TV screen for the telescopes. There is a dark hole near the center of the M5 cover that is used as a reference for aligning the laser beam. If the laser beam doesn't have a well-defined bright or dark spot in the center, you can adjust the focus by opening the "IRIS" gui and clicking [DEC] or [INC] to change the IRIS setting. If you need to, you can click [BEAM] on the IRIS gui; this will close the IRIS and then re-open it to a pre-defined "Beam" position.
  • If the laser beam is far away from the center spot, you can adjust before changing the POP. For example, if you are working on E1, then select "E1M10" from the menu on the "Pico 2" GUI. Then click [MOVE] on the Pico 2 GUI. This will bring up a "PICO E1M10" GUI. Click UP/DOWN/LEFT/RIGHT to center the laser beam on the M10AL spot. (Select and move the correct M10 mirror for the telescope you are aligning.)
  • Now that the laser is aligned, it is time to change the POPs.
  • Look at the POP Overview to check which POP is currently in (the uppermost POP listed as "--"). You will need to remember this later when moving the periscope.
  • On the POP GUI, select the telescope and PoP mirror that needs to go in. For instance, if you need to put in E1-POP2, then select "E1-2" from the menu on the PoPS GUI. Then click [IN] on the PoPS GUI. If you look at the POP Overview, you should see the status of E1-POP2 change to "N" to indicate that it is moving in. Wait for the POP mirror to finish moving in and the status to change to "--" on the Overview.
  • If there are any POPs in that are in front of the selected POP mirror, then those mirrors will need to be moved out. For instance, in the above example, if E1-POP1 is already in, then it will need to be moved out because it will be blocking E1-POP2. To move E1-POP1 out, select "E1-1" from the menu on the PoPS GUI. Then click [OUT] on the PoPS GUI. If you look at the POP Overview, you should see the status of E1-POP1 change to "N" to indicate that it is moving out. Wait for the POP mirror to finish moving out and the status to change to "|" on the Overview. Do this for each POP mirror that is in the way of the selected POP.
  • Note that POP5 = END is ALWAYS IN and listed as "--". It is the last POP mirror in the lab and will never be in the way of any other mirror.
  • During the process above, the laser beam will have disappeared from the "M10AL" screen. It is now time to align the periscope to the new POP mirror.
  • Find the PoP Change Cheat Sheet that is kept between the speaker and VEGA computer on the left side of the monitors. Find the correct telescope chart and the PoP change combination that you are making, eg. E1 PoP 1->2. Take the value in the "From" row and "To" column, (-2.06 in this case) and enter that value in the space with the 1.0 in the proper Periscope for the scope you are changing. If this sheet is unavailable, follow the next paragraph, otherwise skip it.
  • Find the "POP Change Short Procedure" printout that is usually attached to the clipboard in the control room. You will need to look at the table for the "Lower periscope mirror azimuth positions [mm]" on the first page. Compute the difference in the periscope positions from the POP that used to be in to the POP that is currently in. For the above example, changing from "E1-POP1" to "E1-POP2," you will need to move the E1 periscope from 9.01 to 6.945, for a change of -2.06.
  • To move the periscope, click the red button for the telescope on the PERI1 or PERI2 GUI. The power button will turn from red to green to indicate that it is on. The current offset of the periscope is listed in the gray area to the right of the [VEL] button. Note that this offset will be "0.000" if the power on the periscope box in the lab has been turned off and on recently, otherwise there could be some offset listed there. There is a white space between the [-] and [+] buttons where you can enter the amount you need to move the periscope. The values in the PoP Change Cheat Sheet are quite accurate and entering them into the Periscope gui and moving that amount will get you very close each time. The periscope status will change to <---> while it is moving. Wait for it to finish moving (and keep an eye on the M10AL TV screen to look for light). The periscope status will change back to a number when it is finished moving (the status will be the number you moved plus any offset that was originally recorded by the periscope... it's the difference between the original position and the current status that matters). When you see the spot or at least some light, you can change the step size to a small number like "0.05" and click the [-] or [+] button to continue moving the spot in small steps. Wait for the periscope to finish moving between each step. When you have the spot in view, change the step size to an even smaller number like "0.005" and use the [-] and [+] buttons to center as close as possible.
  • When finished, turn the periscope power off by clicking the green power button for the telescope on the PERI1 or PERI2 GUI. The power button for that periscope will turn back to red.
  • You will need to finish the alignment by adjusting the M10 mirror. To do this select, for example, "E1M10" from the menu on the PICO 2 GUI (make sure you choose the M10 for the correct telescope). Then click "MOVE" on the Pico 2 GUI. This will bring up a "PICO E1M10" GUI. Click UP/DOWN/LEFT/RIGHT to center the laser beam on the M10AL spot. Some guis will require increasing the step size to move the spot at a reasonable rate. Many spots will zig-zag as they move, but will continue in the right direction. After you finish, click [OK] on the "PICO E1M10" alignment gui to close the window.
  • Close the shutter for that telescope beam using the SHUTTERS GUI.
  • Open the M5 cover for the telescope. Click ACQ or FIND on the telescope GUI to send the TV back to the window you desire.
  • Repeat the above procedure for each telescope that needs a POP change.
  • After you finish, make sure you change the laser filter back to "ND 3.0". Close the "Laser" shutter on the SHUTTERS GUI. Also make sure that the M5 covers are open on all telescopes.
  • Check the POP Overview to make sure all the POPs are correct. Go to Configuration Tab on Cosmic Debris and click "GET" to get the new configuration. Make sure the POPs update correctly. If the POPs and the ref cart are correct, then click "SEND" to send the new configuration to OPLE and the other servers. You can also check to make sure the configuration is correct on the OPLE GUI Configuration Tab.
  • If you are staying on the same star, go the OPLE GUI and click the [STOP] button at the bottom of the Control Tab. Then click [START] at the bottom of the same tab to restart astrolib in order to send carts to the correct positions for the updated POPs.
  • Make sure the periscope powers are off (red). Close the PER1, PERI2, PICO2, and PoPs GUIs.
  • Update the new POPs on the white board.

Updated 10-22-2016

Troubleshooting

1. A pop mirror does not move IN or OUT

When a pop mirror does not move IN or OUT when clicking on the gui, the most likely solution is to change a blown fuse at the pop mirror. Below is the procedure on how to change a blown fuse.

Tools needed:
Philips screwdriver, A small Allen key, Meter to measure resistance, Tweezers, Replacement fuse (see below at step 1. 4.)

  1. 1. Check the small “POP Mirror Control Box” next to the troubled pop underneath the delay line rails. POP-trouble 1.png (Fig. POP-trouble 1)

  1. 2. Slide the CONTROL switch on the box to “Local”. The red “Timeout” led will most likely turn ON and may be one of the green ones (Fwd or Rev limit) indicating the current position of the pop. POP-trouble 2.png(Fig. POP-trouble 2) Try to use the RESET button: push it with a small enough Allen key. Slide the command switches according to the chart on the box to command the mirror to the other position. Wait about a minute. If the other green led will come on indicating that the mirror moved, then you are done. It was a simple timeout. Otherwise take the next step.


  1. 3. Disconnect all the cables; take the box off the velcro and to a convenient place. Unscrew the four Philips screws at the corners and slowly take off the lid. Do not disconnect the wires inside the box. On the backside of the lid you will see two fuses. Set the meter to beeping mode and check the contact between the two ends of each fuse. ( Fig. POP-trouble 3-4 )
    A beeping fuse is good. You have to replace the fuse, which does not beep. It could be that both of them were blown. POP-trouble 3-4.png

  1. 4. This is the type of fuse that you have to use. Look for the cardboard box labeled “Pop Control Box Fuses” in the BC lab; spare fuses are in there. POP-trouble 5-6.png( Fig. POP-trouble 5-6 )


  1. 5. Use tweezers to take out the blown fuse from its holder, and insert a new one. Put the lid back; be mindful of the right orientation with respect to the connectors. (Look at the first picture.) Secure the lid with the screws, and attach the all the connectors.

  1. 6. Test the box in “Local” mode. Slide the command switches to FWD: 0 1 0 or REV: 0 1 1 to get the mirror moving and hit the Enter button. Wait until the green limit switch led lights up. When the mirror goes against the hard stops, a clinking sound can be heard from the pop box.

  1. 7. If the mirror still does not move as expected further investigation is needed, and probably the pop box has to be opened after letting air into the pipe. If the mirror moved as expected go to the next step.

  1. 8. Slide the CONTROL switch to “Remote”. Make sure you did not leave anything on the rail.